# Symmetric Linear Bandits with Hidden Symmetry

### v1

@@ C.2 Adapting to Separating Constant $\varepsilon_{0}$
The algorithm described in 3 is a direct application of the algorithm proposed in [17], designed for adapting to misspecification errors in linear bandits. Particularly, the algorithm in [17] can adapt to unknown misspecification errors and achieve a regret bound of $\tilde{O}(d_{0}\sqrt{T}+\epsilon_{\mathrm{mis}}T)$, where $\epsilon_{\mathrm{mis}}$ is the misspecification error. At a high level, our algorithm exploring $\sqrt{T}$ rounds using exploratory distribution, which ensures that the misspecification error $\epsilon_{\mathrm{mis}}$ of the chosen subspace $\hat{m}$ is at most $T^{-1/4}$. Therefore, we can run multiple linear bandit algorithms using different levels of misspecification error. Particularly, we use a collection of $K=\lfloor\log(T)\rfloor$ base algorithms, where a base algorithm $k\in[K]$ is a linear bandit algorithm with misspecified level $\varepsilon_{k}=2^{-k}$. Note that the base algorithm $K$ has the same order of regret $\sqrt{T}$ as a well-specified model. Therefore, in exploitation phase, one can guarantee that in the case of well-separated partitions where $\epsilon_{\text{mis}}=0$, the algorithm can achieve a regret of $d_{0}\sqrt{T}$, while in the general case, the regret caused by misspecification error is at most $T^{3/4}\sqrt{d_{0}}$.
@@
See 19
@@
    If $\varepsilon_{0}>T^{-1/4}$, then by Lemma 17, $\theta_{\star}\in\widehat{m}$, that is, $\epsilon_{\mathrm{mis}}=0$;
@@
Appendix D Further Discussions
D.1 Examples on Hidden Subgroups with Subexponential-Size
The model selection procedure in the exploration phase of Algorithms 1 and 2 requires finding the best subspace in the pool $m\in\mathcal{M}$ with respect to least square errors. In the worst case, the algorithm needs to solve $M$ linear regression approaches. Although this step could be sped up using parallel computing, it is still important to exploit certain possible structures in the collection of subspaces $\mathcal{M}$, or equivalently, structures in the collection of partitions $\mathcal{Q}_{d,\leq d_{0}}$, to enable efficient computation. However, we argue that the exact computation of the best subspace $\hat{m}$ as in (5) is an NP-hard problem for many important cases of unknown partitions (or unknown symmetry).
@@ D.1 Examples on Hidden Subgroups with Subexponential-Size
$\mathcal{NC}_{d,\leq d_{0}}\subset\mathcal{Q}_{d,\leq d_{0}}$. One special case of non-crossing partition is interval partition, which has identical structure as sparsity.
@@
We denote $\mathcal{I}_{d}$ as the collection of all interval partition of $d$, we have that $\mathcal{I}_{d}\subset\mathcal{NC}_{d}\subset\mathcal{P}_{d}$.
@@
D.2 Extended Related Work
@@
Sparse linear bandits. As we will explain in Section 4.2, sparsity is equivalent to a subset symmetry structures, and thus, can be seen as a special case of our setting. As such, we first review the literature of sparsity. Sparse linear bandits were first investigated in [3], where the authors achieve a regret of $\tilde{O}(\sqrt{dsT})$, with $\tilde{O}$ disregarding the logarithmic factor, and $s$ representing the sparsity level, and $T$ is the time horizon. Without additional assumptions on the arm set and feature distribution, the lower bound for regret in the sparsity case is $\Omega(\sqrt{dsT})$ [24]. Consequently, the contextual setting has recently gained popularity in the sparsity literature, where additional assumptions are made regarding the context distribution and set of arms. With this assumption, it can be shown that one can achieve regret of the form $\tilde{O}(\tau s\sqrt{T})$, where $\tau$ is a problem-dependent constant that may have a complex form and varies from paper to paper [23, 32]. Apart from the contextual assumption, to avoid polynomial dependence on $T$ in the regret bound, assumptions are required for the set of arms [25, 11, 21]. Recently, [21] offers a unified perspective on the assumption regarding the set of arms by assuming the existence of an exploratory distribution on the set of arms. With this assumption, the authors propose an Explore then Commit style strategy that achieves $\tilde{O}(s^{\frac{2}{3}}T^{\frac{2}{3}})$, nearly matching the lower bound $\Omega(s^{\frac{1}{3}}T^{\frac{2}{3}})$ in the poor data regime [21]. As the sparsity structure can be reduced to a subset of the symmetry structure, all the lower bounds for sparse problems apply to (unknown) symmetric problems.
@@
Symmetry in online learning. The notion of symmetry in Markov Decision Making dates back to works such as [20, 36]. Generally, the reward function and probability transition are preserved under an action of a group on the state-action space. Exploiting known symmetry has been shown to help achieve better performance empirically [39, 38] or tighter regret bounds theoretically [37]. However, all these works requires knowledge of symmetry group, while setting consider unknown subgroup which may be considerably harder. Unknown symmetry on the context or state space has been studied by few authors, with the term context-lumpable bandit [26], meaning that the set of contexts can be partitioned into classes of similar contexts. It is important to note that the symmetry group acts differently on the context space and the action space. As we shall explain in detail in Section 3, while one can achieve a reduction in terms of regret in the case of unknown symmetry acting on context spaces [26], this is not the case when the symmetry group acts on the action space. Particularly, we show that without any extra information on the partition structure of similar classes of arms, no algorithm can achieve any reduction in terms of regret. The work closest to ours is [35], where the authors consider the setting of a $K$-armed bandit, where the set of arms can be partitioned into groups with similar mean rewards, such that each group has at least $q>2$ arms. With the constrained partition, the instance-dependent regret bounds are shown asymptotically to be of order $O\left((\nicefrac{{K}}{{q}})\log T\right)$. Comparing to [35], we study the setting of stochastic linear bandits with similar arms, in which the (unknown) symmetry and linearity structure may intertwine, making the problem more sophisticated. We also impose different constraints on the way one partitions the set of arms, which is more natural in the setting of linear bandits with infinite arms.

### v2

@@ C.2 Adapting to Separating Constant $\varepsilon_{0}$
The algorithm described in 4 is a direct application of the algorithm proposed in [19], designed for adapting to misspecification errors in linear bandits. Particularly, the algorithm in [19] can adapt to unknown misspecification errors and achieve a regret bound of $\tilde{O}(d_{0}\sqrt{T}+\epsilon_{\mathrm{mis}}T)$, where $\epsilon_{\mathrm{mis}}$ is the misspecification error. At a high level, our algorithm exploring $\sqrt{T}$ rounds using exploratory distribution, which ensures that the misspecification error $\epsilon_{\mathrm{mis}}$ of the chosen subspace $\hat{m}$ is at most $T^{-1/4}$. Therefore, we can run multiple linear bandit algorithms using different levels of misspecification error. Particularly, we use a collection of $K=\lfloor\log(T)\rfloor$ base algorithms, where a base algorithm $k\in[K]$ is a linear bandit algorithm with misspecified level $\varepsilon_{k}=2^{-k}$. Note that the base algorithm $K$ has the same order of regret $\sqrt{T}$ as a well-specified model. Therefore, in exploitation phase, one can guarantee that in the case of well-separated partitions where $\epsilon_{\text{mis}}=0$, the algorithm can achieve a regret of $d_{0}\sqrt{T}$, while in the general case, the regret caused by misspecification error is at most $T^{3/4}\sqrt{d_{0}}$.
@@
Corollary 23.
Suppose the Assumptions 5, 6 hold. Then, there exists an algorithm which achieves regret bound as follows:
  • (i)
  • (ii)
    [Non-well-separated partitions] If $\varepsilon_{0}<T^{-1/4}$, then $\mathbf{R}_{T}=\tilde{O}(d_{0}\sqrt{T}+T^{\frac{3}{4}}\sqrt{d_{0}})$.
@@
    If $\varepsilon_{0}>T^{-1/4}$, then by Theorem 18, $\theta_{\star}\in\widehat{m}$, that is, $\epsilon_{\mathrm{mis}}=0$;
@@
Appendix D On Collections of Partitions with Subexponential-Size
D.1 Important Classes of Partitions with Subexponential-Size
In this section, we discuss several important classes of partitions which satisfy Assumption 5.
@@ D.1 Important Classes of Partitions with Subexponential-Size
$\mathcal{NC}_{d,\leq d_{0}}\subset\mathcal{Q}_{d,\leq d_{0}}$.
Another important class of pattern-avoidence partitions is nonnesting partition [14].
Definition 25 (Non-Nesting Partition).
Let $[d]$ admits a cylic order as $1<2<...<d$, and $d<1$. A non-crossing partition of $[d]$ is a partition such that for if $i,\;j$ in one block and $p,\;q$ in one block, then they are not arranged in the order $i<p<q<j$.
Similarly, we denote $\mathcal{NN}_{d}$, $\mathcal{NN}_{d,k}$, $\mathcal{NN}_{d,\leq k}$ as the set of all non-crossing partition of $[d]$, the set of all partition of $[d]$ with $k$ classes, and the set of all partition of $[d]$ with at most $k$ classes. There is bijections between class of $\mathcal{NN}_{d}$ and $\mathcal{NC}_{d}$, non-nesting partitions also satisfy the sub-exponential constraint (Assumption 5).
One special case of both non-crossing partitions and non-nesting partitions is interval partition, which has identical structure as sparsity.
@@
We denote $\mathcal{I}_{d}$ as the collection of all interval partition of $d$, we have that $\mathcal{I}_{d}\subset\mathcal{NC}_{d}\subset\mathcal{P}_{d}$, and $\mathcal{I}_{d}\subset\mathcal{NN}_{d}\subset\mathcal{P}_{d}$.
@@
D.2 Practical Examples of Partitions with Subexponential-Size
General hidden symmetries.
Examples of hidden symmetry in reinforcement learning tasks can be found in robotic control [32, 3], where robot is initially designed symmetrical, but part of symmetry is destroyed by mechanical imperfection. Further examples of hidden symmetry can be also found in the literature on multi-agent reinforcement learning with a large number of agents. To avoid the curse of dimensionality, researchers often rely on the assumption of the existence of homogeneous agents [13, 34]. In the extreme case where all agents are homogeneous, such as in mean-field games, sample complexity becomes independent of the number of agents [13]. However, in practice, agents can be clustered into different types [34], and this information may not be known in advance to the learner (here symmetry occurs between different agents from the same type).
Non-crossing partitions.
Sub-exponential size naturally appears when there is a hierarchical structure on the set $[d]$, and the partitioning needs to respect this hierarchical structure. Particularly, let $T(d,d_{0})$ be the set of ordered trees with $(d+1)$ nodes and $d_{0}$ internal nodes (i.e., nodes that are not the leaves). A partition that respects an ordered tree groups the children of the same node into a single equivalence class (for example, see Figure 1). It is shown in [15] that the cardinality of the set of partitions that respect ordered trees in $T(d,d_{0})$ is sub-exponential. More precisely, it is $O(d^{2d_{0}})$. Furthermore, there is a natural bijection between partitions that respect ordered trees in $T(d,d_{0})$ and the set of non-crossing partitions $\mathcal{NC}_{d,d_{0}}$ [15].
Recall the subcontractor example in the introduction. Here, after the company hires subcontractors $\{1,4,6\}$ to do the job, these subcontractors further break down the tasks into smaller subtasks and hire additional subcontractors $\{2,3\}$, $\{5\}$, and $\{7,8,9\}$, respectively, to execute the subtasks.
Non-nesting partitions.
Besides non-crossing partitions, another sub-exponential-size class of partitions with practical relevance is non-nesting partitions. Consider the resource allocation task where there are $d$ upcoming tasks and $d_{0}$ machines. The job of the designer is to allocate these tasks to each machine.
Now, assume each task will appear in time $t_{1}<t_{2}<\dots<t_{d}$, but the exact time (the value of $t_{i}$) is unknown to the designer. Moreover, the cost of machine $k\in[d_{0}]$, given a subset of tasks $A_{k}$ (ordered according to execution time), is $c_{k}=t_{\max(A_{k})}-t_{\min(A_{k})}$.
The goal of the designer is to minimize the maximum cost of all machines:
$\min\max_{k\in[d_{0}]}c_{k}.$
To achieve this, the designer should avoid nesting allocations (i.e., searching among non-nesting partitions). In particular, assuming that if tasks at times $t_{i}$ and $t_{j}$ are assigned to machine $k$, (where $t_{i}<t_{j}$), and tasks at times $t_{p}$ and $t_{q}$ are assigned to machine $k^{\prime}$ (where $t_{p}<t_{q}$), then it should not be the case that $t_{i}<t_{p}<t_{q}<t_{j}$. This is because the cost of machine $k$ would be significantly higher than that of machine $k^{\prime}$, and the cost could be reduced by swapping task $t_{q}$ for machine $k$ with task $t_{j}$ for machine $k^{\prime}$.
D.3 Efficient Greedy Algorithm for Specific Classes of Partitions
The model selection procedure in the exploration phase of Algorithms 1 and 2 requires finding the best subspace in the pool $m\in\mathcal{M}$ with respect to least square errors. In the worst case, the algorithm needs to solve $M$ linear regression approaches. While the exact computation of the best subspace $\hat{m}$ as in (4) is an NP-hard problem in general (since it contains interval partition as a subclass), we argue that an greedy algorithm can find the ground-truth subspace $m_{\star}$ in $O(nd^{5})$ time complexity, given sufficient large number of samples.
The pseudo code of the greedy algorithm is given in Figure 4. Given that the set of partitions $\mathcal{Q}_{d}$ is equipped with a lattice structure, in which the finest partition is $(1|2|\dots|d)$ and the coarsest partition is $(1,2,\dots,d)$.
The algorithm starts with the finest partition $\hat{\pi}=(1|2|\dots|d)$. In each iteration, the algorithm finds the finest coarsening of the current partition $\hat{\pi}$. In graph-theoretic terms, it finds all neighbors of $\hat{\pi}$ in the lattice that are coarsenings of $\hat{\pi}$. The operator for finding the finest coarsening of $\hat{\pi}$ is denoted as Coarsen($\hat{\pi}$), and it returns a collection of the finest coarsened partitions. Next, the algorithm finds the partition that minimizes the prediction error $\|Y-\Pi_{S_{m}}(Y)\|_{2}^{2}$ among the current coarsening collection. At the end of each iteration, as the number of classes in $\hat{\pi}$ is reduced by one, the dimension variable is also reduced by $1$. The while loop stops when the dimension equals $d_{0}$.
Since the algorithm only optimizes locally within the current coarsening collection, it exhibits a behavior similar to a greedy algorithm. We note that for non-crossing and non-nesting partitions, the cardinality of the coarsening collection at any level of the lattice is at most $d^{2}$. Therefore, assuming that creating a finest coarsening partition take $O(d)$ operator, and solving least square takes $O(nd^{2})$, the algorithms time complexity is $O(nd^{5})$.
Algorithm 4 Greedily Search within Lattice
1:  Input: Compact representation of $\mathcal{M}$, design matrix $X$, reward vector $Y$.
2:  Initialise $\hat{\pi}=(1|2|3|4...|d)$, $\mathrm{dimension}=d$.
3:  while $\mathrm{dimension}>d_{0}$ do
4:     Collection $=$ Coarsen($\hat{\pi}$).
5:     $\hat{m}=\operatorname*{arg\,min}_{m\in\mathbf{H}(\mathrm{Collection})}\|Y-\Pi_{S_{m}}(Y)\|_{2}^{2}$.
6:     $\hat{\pi}=\mathbf{H}^{-1}(\hat{m})$.
7:     $\mathrm{dimension}=\mathrm{dimension}-1$.
8:  end while
9:  $\hat{\theta}=\operatorname*{arg\,min}_{\theta\in\hat{m}}\|Y-X\theta\|_{2}^{2}$.
10:  Return $\hat{m},\;\hat{\theta}$.
Appendix E Experiment details
We conduct simulations where the entries of $\theta_{\star}$ satisfy non-crossing partition constraints. The set of arms $\mathcal{X}$ is $\sqrt{d}\mathbb{S}^{d-1}$, $\sigma=0.1$, and $(d,d_{0})\in\{(40,4),(80,10),(100,15)\}$. We let exploratory distribution $\nu$ be the uniform distribution on the unit sphere. The ground-truth partition $\pi_{\mathcal{G}}$ and $\theta_{\star}$ are randomized before each simulation.
To run ESTC-Lasso algorithm [23], we introduce an auxiliary sparse vector $\varphi$ corresponding to $\theta_{\star}$, whose entries are defined as $\varphi_{i}=\theta_{i+1}-\theta_{i}$, and $\varphi_{d}=\theta_{d}$. We apply Lasso regression for $\varphi_{\star}$, get the estimate $\hat{\varphi}$, then convert back to $\hat{\theta}$ using the map that transforms sparse vector to interval-partition vector (inversion of the map we defined above).
Regarding implementing our algorithm, we use greedy Algorithm 4 as introduced in Appendix D.3, to solve the optimisation in equation (3), (4), and its complexity is $O(t_{1}d^{5})$. It is shown in the simulation result (Figure 2, 3, 4), the greedy algorithm achieves small risk error and consequently leads to small regret. Code is available at:
https://github.com/NamTranKekL/Symmetric-Linear-Bandit-with-Hidden-Symmetry.git.
Figure 3: Regret of EMC (Algorithm 1) and of ESTC proposed in [23], in cases of sparsity, non-crossing partitions, and non-nesting partitions, with $d=40,\;d_{0}=4$.
Figure 4: Regret of EMC (Algorithm 1) and of ESTC proposed in [23], in cases of sparsity, non-crossing partitions, and non-nesting partitions, with $d=80,\;d_{0}=10$.
Appendix F Extended Related Work
@@
Sparse linear bandits. As we will explain in Section 4.2, sparsity is equivalent to a subset symmetry structures, and thus, can be seen as a special case of our setting. As such, we first review the literature of sparsity. Sparse linear bandits were first investigated in [2], where the authors achieve a regret of $\tilde{O}(\sqrt{dsT})$, with $\tilde{O}$ disregarding the logarithmic factor, and $s$ representing the sparsity level, and $T$ is the time horizon. Without additional assumptions on the arm set and feature distribution, the lower bound for regret in the sparsity case is $\Omega(\sqrt{dsT})$ [27]. Consequently, the contextual setting has recently gained popularity in the sparsity literature, where additional assumptions are made regarding the context distribution and set of arms. With this assumption, it can be shown that one can achieve regret of the form $\tilde{O}(\tau s\sqrt{T})$, where $\tau$ is a problem-dependent constant that may have a complex form and varies from paper to paper [26, 36]. Apart from the contextual assumption, to avoid polynomial dependence on $T$ in the regret bound, assumptions are required for the set of arms [28, 11, 23]. Recently, [23] offers a unified perspective on the assumption regarding the set of arms by assuming the existence of an exploratory distribution on the set of arms. With this assumption, the authors propose an Explore then Commit style strategy that achieves $\tilde{O}(s^{\frac{2}{3}}T^{\frac{2}{3}})$, nearly matching the lower bound $\Omega(s^{\frac{2}{3}}T^{\frac{2}{3}})$ in the poor data regime [24]. As the sparsity structure can be reduced to a subset of the symmetry structure, all the lower bounds for sparse problems apply to (unknown) symmetric problems.
@@
Symmetry in online learning. The notion of symmetry in Markov Decision Making dates back to works such as [22, 40]. Generally, the reward function and probability transition are preserved under an action of a group on the state-action space. Exploiting known symmetry has been shown to help achieve better performance empirically [43, 42] or tighter regret bounds theoretically [41]. However, all these works requires knowledge of symmetry group, while setting consider unknown subgroup which may be considerably harder. Unknown symmetry on the context or state space has been studied by few authors, with the term context-lumpable bandit [29], meaning that the set of contexts can be partitioned into classes of similar contexts. It is important to note that the symmetry group acts differently on the context space and the action space. As we shall explain in detail in Section 3, while one can achieve a reduction in terms of regret in the case of unknown symmetry acting on context spaces [29], this is not the case when the symmetry group acts on the action space. Particularly, we show that without any extra information on the partition structure of similar classes of arms, no algorithm can achieve any reduction in terms of regret.
The work closest to ours is [39], where the authors consider the setting of a $K$-armed bandit, where the set of arms can be partitioned into groups with similar mean rewards, such that each group has at least $q>2$ arms. With the constrained partition, the instance-dependent regret bounds are shown asymptotically to be of order $O\left((\nicefrac{{K}}{{q}})\log T\right)$. Comparing to [39], we study the setting of stochastic linear bandits with similar arms, in which the (unknown) symmetry and linearity structure may intertwine, making the problem more sophisticated. We also impose different constraints on the way one partitions the set of arms, which is more natural in the setting of linear bandits with infinite arms. As a result, we argue that the technique used in [39] cannot be applied in our setting. To clarify this, let us first review the algorithmic technique from [39]: The algorithm in [39] assumes there is an equivalence among the parameters $\theta$, and that the set of arms $\mathcal{X}$ is a simplex. At each round $t$, given an estimation $\hat{\theta}_{t}$, the algorithm maintains a sorted list of indices in $[d]$ that follows the ascending order of the magnitude of $\hat{\theta}_{i}$. The algorithm then uses the sorted list $(\hat{\theta}_{i})_{i\in[d]}$ to choose arm $x$ accordingly. The key assumption here is that, since the set of arms $\mathcal{X}$ is a simplex, we can estimate each $\theta_{i}$ independently. This implies that the list $(\hat{\theta}_{i})_{i\in[d]}$ should respect the true order of the list $(\theta_{i})_{i\in[d]}$ when there are a sufficiently large number of samples. Unfortunately, this is typically not the case in linear bandits where $\mathcal{X}$ has a more general shape. In linear bandits, there can be correlations between the estimates $\hat{\theta}_{i}$ and $\hat{\theta}_{j}$ for any $i,j\in[d]$. Hence, one should not expect that $(\hat{\theta}_{i})_{i\in[d]}$ will maintain the same order as $(\theta_{i})_{i\in[d]}$. In other words, the correlations among the estimates $\{\hat{\theta}_{1},\ldots,\hat{\theta}_{d}\}$ may destroy the original order in $\{\theta_{1},\ldots,\theta_{d}\}$. In fact, we can only guarantee the risk error of estimation $\theta$, i.e., $\|\hat{\theta}-\theta_{\star}\|$ is small, but not necessarily the order of the indices in $\theta$. Therefore, the technique used in [39] cannot be directly applied to linear bandits in its current form.